OpenStudy (anonymous):

What is the surface area of the square pyramid? A. 2242 in2 B. 3838 in2 C. 4636 in2 D. 30,324 in2

OpenStudy (anonymous):

ill give medel

OpenStudy (welshfella):

required area = area pf the base + 4 * area of each triangular side

OpenStudy (welshfella):

area of the base = 38^2 because it is a square and the area of a triangle = 1/2 * base * height - just plug in the given measures into this formula

OpenStudy (anonymous):

i dont get it :/

OpenStudy (welshfella):

which part don't you get?

OpenStudy (anonymous):

how do i do the math

OpenStudy (welshfella):

well you can first work out 38^2 = 38 * 38 - use ur calculator

OpenStudy (anonymous):

it's = 1444

OpenStudy (welshfella):

right now the base of each triangle = same as the side of the base = 38 in and the height is given as 42 in so the area of one triangle = 1/2 * 38 * 42

OpenStudy (anonymous):

798.5

OpenStudy (aaronandyson):

once you do that add 1444 and you get your final ans :)

OpenStudy (aaronandyson):

OMG!! ITS 1596 NOT 97 SO SORRY

OpenStudy (aaronandyson):

38*42 = 1596

OpenStudy (welshfella):

you have also to multiply 798 by since there are 4 triangles in the pyramid

OpenStudy (welshfella):

* 798 by 4

OpenStudy (welshfella):

after multiplying 798 by 4 add the result to 1444 and you have the answer